# Nonlinear Separation Theorems for Co-Radiant Sets and Optimality Conditions for Approximate and Proper Approximate Solutions in Vector Optimization

### v2

@@ 4. Separation Theorems for Co-Radiant Sets
In this section, we derive necessary and sufficient conditions for approximate solutions via sublinear scalar optimization problems. We begin establishing some notation and terminology in order to introduce the notion of $\varepsilon$-efficient point. Let $X$ be a normed space and $C\subset X$ a proper pointed convex cone. The cone $C$ induces a partial order $\leq$ on $X$ defined as follows: for $x,y\in X$, $y\leq x$ if and only if $x-y\in C$. In this context, given a non-empty set $A\subset X$, a point $x_{0}\in A$ is said to be an efficient point of $A$, denoted as $x_{0}\in\mbox{E}(A,C)$, if $A\cap(-C\setminus\{0_{X}\})=\emptyset$. Let introduce the concept of $\varepsilon$-efficiency by adapting [8, Definition 3.2] to our context. This notion is based on replacing the ordering cone with a proper pointed co-radiant set $\mathcal{C}$ that approximates it. Recall that we define $\mathcal{C}(\varepsilon)=\varepsilon\mathcal{C}$ for every $\varepsilon>0$.
@@
Let us check that ${{\mathcal{K}}}_{f,\alpha}\not=\emptyset$. Indeed, taking some $x_{0}\in S_{X}$ such that $f(x_{0})>\alpha$, we get that $f(x_{0})-\alpha\|x_{0}\|>0$. Consequently, there exists $\mu>0$ such that $f(\mu x_{0})-\alpha\|\mu x_{0}\|>1$, whence $\mu x_{0}\in{{\mathcal{K}}}_{f,\alpha}$. Then, ${{\mathcal{K}}}_{f,\alpha}$ is proper because $0_{X}\not\in{{\mathcal{K}}}_{f,\alpha}$. The fact that ${{\mathcal{K}}}_{f,\alpha}$ is a co-radiant subset is straightforward. To check that ${{\mathcal{K}}}$ is pointed we assume that $x\in{{\mathcal{K}}}_{f,\alpha}\cap(-{{\mathcal{K}}}_{f,\alpha})$. Then $f(x)+\alpha\|x\|<-1<1<f(x)-\alpha\|x\|$, that implies $\alpha<0$ (because $x\not=0_{X}$), a contradiction. ∎
@@
Taking into account that $-{{\mathcal{K}}}_{f,\alpha}(\varepsilon)=\{x\in X\colon p_{f,\alpha}(x)<-\lambda\}$ the following chain of equivalences becomes immediate. Indeed, $x_{0}\in\mbox{AE}(A,K_{f,\alpha},\lambda)$Leftrightarrow $(A-x_{0})\cap(-{{\mathcal{K}}}_{f,\alpha}(\lambda))=\varnothing$Leftrightarrow $(A-x_{0})\subset\{x\in X\colon p_{f,\alpha}(x^{\prime}-x_{0})\geq-\lambda\}$Leftrightarrow $p_{f,\alpha}(x^{\prime}-x_{0})\geq-\lambda\ \forall x^{\prime}\in A.$
@@
Let $X$ be a normed space and let ${{\mathcal{C}}}\subset X$ be a co-radiant subset such that $d_{{{\mathcal{C}}}}>0$. If $\mbox{cone}({{\mathcal{C}}})$ has a bounded base, then there exists $\lambda>0$ such that ${{\mathcal{C}}}^{a*}_{\lambda}(\varepsilon)\not=\emptyset$ and ${{\mathcal{C}}}^{a\#}_{\frac{\lambda}{2}}(\varepsilon)\not=\emptyset$, $\forall\varepsilon>0$.
@@
Let ${{\mathcal{C}}}$ be a co-radiant set such that $d_{{{\mathcal{C}}}}>0$ and $C:=\mbox{cone}({{\mathcal{C}}})$ has a bounded base. By [3, Theorem 1.1], there exists $f\in S_{X^{*}}$ such that $\gamma:=\inf_{C\cap S_{X}}f>0$. Then $f(\frac{c}{\|c\|})\geq\gamma$, $\forall c\in{{\mathcal{C}}}$Rightarrow $f(c)-\gamma\|c\|\geq 0$, $\forall c\in{{\mathcal{C}}}$. Now, fix arbitrary $\varepsilon>0$ and $x\in{{\mathcal{C}}}(\varepsilon)$. Then $f(x)-\frac{\gamma}{2}\|x\|=f(x)-\gamma\|x\|+\frac{\gamma}{2}\|x\|\geq\frac{\gamma}{2}\|x\|\geq\frac{\gamma\varepsilon d_{{{\mathcal{C}}}}}{2}$. Then, taking $\lambda:=\frac{\gamma\varepsilon d_{{{\mathcal{C}}}}}{2}$ we have that $(f,\frac{\gamma}{2})\in{{\mathcal{C}}}^{a*}_{\lambda}(\varepsilon)$ and $(f,\frac{\gamma}{2})\in{{\mathcal{C}}}^{a\#}_{\frac{\lambda}{2}}(\varepsilon)$. ∎
@@
    If $10<\varepsilon<$ 1, $\alpha\in(\alpha_{1},\alpha_{2})$, and $x_{0}\in AE(A,{{\mathcal{C}}}_{\delta}(\eta),\varepsilon)$, then:
@@
    $\left[x_{0}+\left\{x\in X:p_{f,\alpha}(x)<0\right\}\right]\cap A\subset AE(A,{{\mathcal{C}}},\varepsilon).$ (12)
@@
The condition ${{\mathcal{C}}}\cap nS_{X}\not=\emptyset$ implies that $\overline{\mbox{co}}({{\mathcal{C}}})\cap{{\mathcal{C}}}_{\delta}\cap\mbox{cone}({{\mathcal{C}}}_{\delta}\cap nS_{X})\not=\emptyset$. Then, we apply Theorem 4.9 and we have $f\in S_{X^{*}}$, $0<\alpha_{1}<\alpha_{2}$ such that $\lambda:=\inf\{f(x^{\prime})-\alpha\|x^{\prime}\|:x^{\prime}\in{{\mathcal{C}}},\,\alpha\in(\alpha_{1},\alpha_{2})\}>0$, and there exists $0<\eta\leq 1$ such that for $f(y)-\alpha\|y\|<\lambda<f(x^{\prime})-\alpha\|x^{\prime}\|$, for every $\alpha_{1}<\alpha<\alpha_{2}$, $y\in X\setminus\mbox{int}({{\mathcal{C}}}_{\delta}(\eta))$ and $x^{\prime}\in\overline{\mbox{co}}({{\mathcal{C}}})$. To check $\alpha_{2}\leq 1$, consider arbitrary $\alpha_{1}<\alpha<\alpha_{2}$ and $x^{\prime}\in{{\mathcal{C}}}$. Then $f(x^{\prime})>\alpha\|x^{\prime}\|$, which implies that $f(\frac{x^{\prime}}{\|x^{\prime}\|})>\alpha$, $\forall x^{\prime}\in{{\mathcal{C}}}$. As a consequence, $f\in{{\mathcal{C}}}^{\#}$ and $1>\alpha$. As $\alpha_{1}<\alpha<\alpha_{2}$ was arbitrarily taken, the inequality is proved. In particular, we also have that $(f,\alpha)\in{{\mathcal{C}}}^{a\#}_{\lambda\varepsilon}(\varepsilon)$, $\forall\varepsilon>0$. On the other hand, it is clear that $p_{f,\alpha}(x^{\prime})<-\lambda<p_{f,\alpha}(y)$, for every $\alpha_{1}<\alpha<\alpha_{2}$, $y\in Y\setminus\mbox{int}(-{{\mathcal{C}}}_{\delta}(\eta))$ and $x^{\prime}\in\overline{\mbox{co}}(-{{\mathcal{C}}})$. Let us fix some $10<\varepsilon<$ 1. Then, $p_{f,\alpha}(y)>-\lambda\varepsilon$, for every $y\in Y\setminus\mbox{int}(-{{\mathcal{C}}}_{\delta}(\eta\varepsilon))$ and $\alpha_{1}<\alpha<\alpha_{2}$. Finally, since $x_{0}\in\mbox{AE}(A,{{\mathcal{C}}}_{\delta}(\eta),\varepsilon)$, we have the inclusion $A\subset X\setminus(x_{0}-{{\mathcal{C}}}_{\delta}(\eta\varepsilon))$, which implies that $p_{f,\alpha}(x-x_{0})>-\lambda\varepsilon$, for $x\in A$ proving (11). Finally, since $(f,\alpha)\in{{\mathcal{C}}}^{a*}_{\lambda\varepsilon}(\varepsilon)$, we can apply Theorem 5.5 and get
@@
Let us introduce another notion to be used in our last remark. Fix $g:X\rightarrow\mathbb{R}$, $A\subset X$, and $A\not=\emptyset$. For every $\varepsilon>0$, it is defined the set of $\varepsilon$-approximate solutions of $\mbox{Min}\{g(x)\colon x\in A\}$ as $\text{AMin}(g,A,\varepsilon):=\{x_{0}\in A\colon g(x_{0})-\varepsilon\leq g(x),\,\forall x\in A\}.$ It is easy to check that $\cap_{\varepsilon>0}\text{AMin}(g,A,\varepsilon)=\mbox{Min}\{g(x)\colon x\in A\}$. The following formula
@@
In [8], some necessary and sufficient conditions for $\varepsilon$-efficiency were obtained via $\varepsilon$-approximate solutions of related scalar optimization problems. For necessary conditions, the authors had to assume that ${{\mathcal{C}}}$ is star-shaped with kern$({{\mathcal{C}}})$ solid and, for sufficient conditions, they had to use an extra strictly local monotone function. In our approach, we find it more convenient to consider suitable functions $g$ that are sublinear; in particular, the role of $g$ here is played by $p_{f,\alpha}$. This choice allows us not only to avoid the restrictions on ${{\mathcal{C}}}$, but consider a bigger set of solutions for the scalarizations.

### v3

@@ 4. Separation Theorems for Co-Radiant Sets
In this section, we derive necessary and sufficient conditions for approximate solutions via sublinear scalar optimization problems. We begin establishing some notation and terminology in order to introduce the notion of $\varepsilon$-efficient point. Let $X$ be a normed space and $C\subset X$ a proper pointed convex cone. The cone $C$ induces a partial order $\leq$ on $X$ defined as follows: for $\in X, yxy\leq x$ if and only if $x-y\in C$. In this context, given a non-empty set $A\subset X$, a point $x_{0}\in A$ is said to be an efficient point of $A$, denoted as $x_{0}\in\mbox{E}(A,C)$, if $A\cap(-C\setminus\{0_{X}\})=\emptyset$. Let introduce the concept of $\varepsilon$-efficiency by adapting [8, Definition 3.2] to our context. This notion is based on replacing the ordering cone with a proper pointed co-radiant set $\mathcal{C}$ that approximates it. Recall that we define $\mathcal{C}(\varepsilon)=\varepsilon\mathcal{C}$ for every $\varepsilon>0$.
@@
Let us check that ${{\mathcal{K}}}_{f,\alpha}\not$$\emptyset$. Indeed, taking some $x_{0}\in S_{X}$ such that $f(x_{0})>\alpha$, we get that $f(x_{0})-\alpha\|x_{0}\|>0$. Consequently, there exists $\mu>0$ such that $f(\mu x_{0})-\alpha\|\mu x_{0}\|>1$, whence $\mu x_{0}\in{{\mathcal{K}}}_{f,\alpha}$. Then, ${{\mathcal{K}}}_{f,\alpha}$ is proper because $0_{X}\not\in{{\mathcal{K}}}_{f,\alpha}$. The fact that ${{\mathcal{K}}}_{f,\alpha}$ is a co-radiant subset is straightforward. To check that $\mathcal{K}$}} is pointed we assume that $x\in{{\mathcal{K}}}_{f,\alpha}\cap(-{{\mathcal{K}}}_{f,\alpha})$. Then $f(x)+\alpha\|x\|<-1<1<f(x)-\alpha\|x\|$, that implies $\alpha<0$ (because $x\not=0_{X}$), a contradiction. ∎
@@
Taking into account that ${{\mathcal{K}}}_{f,\alpha}(\varepsilon)=\{x\in X\colon p_{f,\alpha}(x)<-\lambda\}$ the following chain of equivalences becomes immediate. Indeed, $x_{0}\in\mbox{AE}(A,K_{f,\alpha},\lambda)$ $\Leftrightarrow$ $(A-x_{0})\cap(-{{\mathcal{K}}}_{f,\alpha}(\lambda))=\varnothing$ $\Leftrightarrow$ $(A-x_{0})\subset\{x\in X\colon p_{f,\alpha}(x^{\prime}-x_{0})\geq-\lambda\}$ $\Leftrightarrow$ $p_{f,\alpha}(x^{\prime}-x_{0})\geq-\lambda\ \forall x^{\prime}\in A.$
@@
Let $X$ be a normed space and let ${{\mathcal{C}}}\subset X$ be a co-radiant subset such that $d_{{{\mathcal{C}}}}>0$. If $\mbox{cone}({{\mathcal{C}}})$ has a bounded base, then there exists $\lambda>0$ such that ${{\mathcal{C}}}^{a*}_{\lambda}(\varepsilon)\not$$\emptyset$ and ${{\mathcal{C}}}^{a\#}_{\frac{\lambda}{2}}(\varepsilon)\not$$\emptyset$, $\forall\varepsilon>0$.
@@
Let $\mathcal{C}$}} be a co-radiant set such that $d_{{{\mathcal{C}}}}>0$ and $C:=\mbox{cone}({{\mathcal{C}}})$ has a bounded base. By [3, Theorem 1.1], there exists $f\in S_{X^{*}}$ such that $\gamma:=\inf_{C\cap S_{X}}f>0$. Then $f(\frac{c}{\|c\|})\geq\gamma$, $\forall c\in{{\mathcal{C}}}$ $\Rightarrow$ $f(c)-\gamma\|c\|\geq 0$, $\forall c\in{{\mathcal{C}}}$. Now, fix arbitrary $\varepsilon>0$ and $x\in{{\mathcal{C}}}(\varepsilon)$. Then $f(x)-\frac{\gamma}{2}\|x\|=f(x)-\gamma\|x\|+\frac{\gamma}{2}\|x\|\geq\frac{\gamma}{2}\|x\|\geq\frac{\gamma\varepsilon d_{{{\mathcal{C}}}}}{2}$. Then, taking $\lambda:=\frac{\gamma\varepsilon d_{{{\mathcal{C}}}}}{2}$ we have that $(f,\frac{\gamma}{2})\in{{\mathcal{C}}}^{a*}_{\lambda}(\varepsilon)$ and $(f,\frac{\gamma}{2})\in{{\mathcal{C}}}^{a\#}_{\frac{\lambda}{2}}(\varepsilon)$. ∎
@@
    If $0<\varepsilon<1$, $\alpha\in(\alpha_{1},\alpha_{2})$, and $x_{0}\in AE(A,{{\mathcal{C}}}_{\delta}(\eta),\varepsilon)$, then:
@@
    $\left[x_{0}+\left\{x\in X:p_{f,\alpha}(x)<0\right\}\right]\cap A\subset AE(A,{{\mathcal{C}}},\varepsilon).$
@@
The condition ${{\mathcal{C}}}\cap nS_{X}\not$$\emptyset$ implies that $\overline{\mbox{co}}({{\mathcal{C}}})\cap{{\mathcal{C}}}_{\delta}\cap\mbox{cone}({{\mathcal{C}}}_{\delta}\cap nS_{X})\not$$\emptyset$. Then, we apply Theorem 4.9 and we have $f\in S_{X^{*}}$, $0<\alpha_{1}<\alpha_{2}$ such that $\lambda:=\inf\{f(x^{\prime})-\alpha\|x^{\prime}\|:x^{\prime}\in{{\mathcal{C}}},\,\alpha\in(\alpha_{1},\alpha_{2})\}>0$, and there exists $0<\eta\leq 1$ such that for $f(y)-\alpha\|y\|<\lambda<f(x^{\prime})-\alpha\|x^{\prime}\|$, for every $\alpha_{1}<\alpha<\alpha_{2}$, $y\in X\setminus\mbox{int}({{\mathcal{C}}}_{\delta}(\eta))$ and $x^{\prime}\in\overline{\mbox{co}}({{\mathcal{C}}})$. To check $\alpha_{2}\leq 1$, consider arbitrary $\alpha_{1}<\alpha<\alpha_{2}$ and $x^{\prime}\in{{\mathcal{C}}}$. Then $f(x^{\prime})>\alpha\|x^{\prime}\|$, which implies that $f(\frac{x^{\prime}}{\|x^{\prime}\|})>\alpha$, $\forall x^{\prime}\in{{\mathcal{C}}}$. As a consequence, $f\in{{\mathcal{C}}}^{\#}$ and $1>\alpha$. As $\alpha_{1}<\alpha<\alpha_{2}$ was arbitrarily taken, the inequality is proved. In particular, we also have that $(f,\alpha)\in{{\mathcal{C}}}^{a\#}_{\lambda\varepsilon}(\varepsilon)$, $\forall\varepsilon>0$. On the other hand, it is clear that $p_{f,\alpha}(x^{\prime})<-\lambda<p_{f,\alpha}(y)$, for every $\alpha_{1}<\alpha<\alpha_{2}$, $y\in Y\setminus\mbox{int}(-{{\mathcal{C}}}_{\delta}(\eta))$ and $x^{\prime}\in\overline{\mbox{co}}(-{{\mathcal{C}}})$. Let us fix some $0<\varepsilon<1$. Then, $p_{f,\alpha}(y)>-\lambda\varepsilon$, for every $y\in Y\setminus\mbox{int}(-{{\mathcal{C}}}_{\delta}(\eta\varepsilon))$ and $\alpha_{1}<\alpha<\alpha_{2}$. Finally, since $x_{0}\in\mbox{AE}(A,{{\mathcal{C}}}_{\delta}(\eta),\varepsilon)$, we have the inclusion $A\subset X\setminus(x_{0}-{{\mathcal{C}}}_{\delta}(\eta\varepsilon))$, which implies that $p_{f,\alpha}(x-x_{0})>-\lambda\varepsilon$, for $x\in A$ proving (10). Finally, since $(f,\alpha)\in{{\mathcal{C}}}^{a*}_{\lambda\varepsilon}(\varepsilon)$, we can apply Theorem 5.5 and get
@@
Let us introduce another notion to be used in our last remark. Fix $g:X\rightarrow\mathbb{R}$, $A\subset X$, and $A\not$$\emptyset$. For every $\varepsilon>0$, it is defined the set of $\varepsilon$-approximate solutions of $\mbox{Min}\{g(x)\colon x\in A\}$ as $\text{AMin}(g,A,\varepsilon):=\{x_{0}\in A\colon g(x_{0})-\varepsilon\leq g(x),\,\forall x\in A\}.$ It is easy to check that $\cap_{\varepsilon>0}\text{AMin}(g,A,\varepsilon)=\mbox{Min}\{g(x)\colon x\in A\}$. The following formula
@@
[…]